OpenStudy (anonymous):

Find the present value, using the present value formula and a calculator. (Round your answer to the nearest cent.) Achieve $225,500 at 8.55% compounded continuously for 8 years, 155 days.

OpenStudy (bahrom7893):

What's the present value formula?

OpenStudy (anonymous):

\[m \left[ (1-(1+r/n)^{-nt})\div(r/n ) \right]\]

OpenStudy (anonymous):

part im confused on is the compounded continuously

OpenStudy (bahrom7893):

For continuous compounding, use the formula with e: A = Pe^(rt)

OpenStudy (anonymous):

so i would have 225,500=Pe^(.0855* 8.42)? How do i solve for P?

OpenStudy (bahrom7893):

Divide both sides by e^(.0855*8.42)

OpenStudy (anonymous):

Thats it?

OpenStudy (bahrom7893):

yup

OpenStudy (bahrom7893):

e is a constant, remember.. e = 2.71...

OpenStudy (anonymous):

right

OpenStudy (calculusfunctions):

Sorry, my server keeps crashing, but The present value of a perpetuity under which you receive amount A annually is A / r where r is the annual interest rate, compounded continuously. You can prove the formula with calculus.

OpenStudy (anonymous):

A is 225,500?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

so just do 225,500/.0855? Thats your answer?

OpenStudy (bahrom7893):

no, do: 225,500/e^(.0855*8.42)

OpenStudy (tkhunny):

First, you shoudl get the right formula. Are you SURE those should be 'n' in the parentheses and in the denominator? Second, just run an experiment. Substitute greater and greater values for 'm'. See if it settles down.